If momentum of a body be increased by 50% then it KE will increased by? 
 

Grade:12th pass

1 Answers

na batau
5 years ago
use relation p={2*m*(KE)}^(1/2)
p1={2*m*(KE-1)}^(1/2)  …..    ….           ….---(1)
p2=(3/2)*(p1)={2*m*(KE-2)}^(1/2)............ …. . .. . . . . ----(2)
 
divide equation (1) and (2) to get
(KE-2)=(9/4)*(KE-1)